OpenStudy (anonymous):

A wheel of radius 1.0 m turns by an angle of 77°. How long is the arc of the displacement, s?

OpenStudy (anonymous):

\[\theta = \frac{ arc }{ radius }\] \[arc = \theta * radius\] Remember Angle Shuld be in Radians